# Long-Lived Circular Rydberg Qubits of Alkaline-Earth Atoms in Optical Tweezers

### 19 Jan 2024 v1

@@ V Lifetime
Finally, we investigate the lifetime of our circular Rydberg qubit in more detail. To this end, we initialize the qubit either in |$\ket{79\mathrm{C}}$ or |$\ket{77\mathrm{C}}$, the latter via a microwave $\pi$-pulse applied after preparing |$\ket{79\mathrm{C}}$. For both scenarios, we hold the atoms in shallow tweezers ($\approx 400\,\mathrm{\upmu W}$) for a variable time $t$ and subsequently perform state-selective field ionization. This allows us to identify black-body induced population transfer into neighboring $n$-manifolds. Representative histograms of the field-ionization signal are depicted in Fig. 5(c) and (d) at $t\approx 630\,\mathrm{\upmu s}$ (indicated by dashed vertical lines) when starting in either of the qubit states. Evidently, decay from the circular state appears to be significantly slower for |$\ket{79\mathrm{C}}$. For a quantitative analysis, we fit the histograms for each hold time with a sum over multiple skewed Gaussians with the individual amplitudes as only free parameters and extract the time-dependent state population $p_{n}$ from the areas under the individual curves. Results are depicted in Fig. 5(a) and (b), and allow for extracting lifetimes $\tau_{\ket{n\mathrm{C}}}$ by fitting the data with a rate model (see Appendix E). Specifically, we obtain $\tau_{\ket{79\mathrm{C}}}=2.55(10)\,\mathrm{ms}$ and $\tau_{\ket{77\mathrm{C}}}=0.53(8)\,\mathrm{ms}$. Comparing these observations to predictions in free space at room temperature ($\tau_{\ket{79\mathrm{C}},\rm{fs}}=303\,\mathrm{\upmu s}$ and $\tau_{\ket{77\mathrm{C}},\rm{fs}}=297\,\mathrm{\upmu s}$), we find an enhancement in lifetime by factors of $8.4$ and $1.8$, respectively.
The long circular state lifetimes are attributed to the presence of the ITO electrodes forming a plate capacitor in the $x$-$y$ plane, which coincides with the orbital plane of the Rydberg electron. The capacitor plates are spaced by $d=1\,\mathrm{cm}$, which is slightly smaller than the half-wavelength of black-body transitions into the neighboring $n$-manifolds $\lambda_{\rm{BB}}^{n\rightarrow n-1}/2=1.06~{}(1.02)\,\mathrm{cm}$ for |$\ket{79\mathrm{C}}$ (|$\ket{77\mathrm{C}}$). This leads to suppression of the most detrimental circularly polarized black-body modes inside the electrode structure and allows us to create and control long-lived circular states without cryogenic cooling. The fact that the capacitor spacing is only slightly larger than the relevant black-body wavelength leads to the large difference in suppression factors we find for the two circular qubit states.
@@ VI Conclusion and Outlook
We have demonstrated creation and trapping of alkaline-earth atoms in very high-$n$ circular Rydberg states by applying radio-frequency driven adiabatic state preparation to a record-breaking principal quantum number of $n=79$. Trapping in a standard Gaussian tweezer beam is possible due to the core polarization of the divalent Rydberg atom. We implemented microwave control of a circular Rydberg atom qubit and characterized qubit coherence via Ramsey and spin-echo spectroscopy, including analysis of trap-induced dephasing. Our results show coherence times up to 278(10) $\upmu$s, limited by residual magnetic field noise in our current setup. Assisted by an in-vacuum suppression capacitor for black-body modes, we find that the produced circular states live for up to 2.5 ms, which to the best of our knowledge is the longest-lived Rydberg atom ever observed in a room temperature environment.
Our work opens the door to quantum simulations with alkaline-earth circular Rydberg states, and provides prospects for novel qubit concepts in gate-based Rydberg quantum computers [13, 14, 17]. Improved magnetic field noise cancellation should allow for reaching qubit coherence in the millisecond range and beyond [17], which provides exciting opportunities to overcome fundamental lifetime limitations in state-of-the-art quantum simulators and quantum computers using low-$l$ Rydberg states by orders of magnitude. The available ionic-core transitions can be exploited in future experiments for optical read-out and laser cooling of circular Rydberg atoms [33, 34, 35]. Narrow-line spectroscopy on the ${}^{+}$ core also enables coherent and local optical manipulation of the circular state qubit mediated by the electrostatic coupling between the two electrons [27].
Next obvious steps include the creation and control of interacting circular Rydberg atoms, scaling the system to large qubit arrays, possibly assisted by rapid Rabi coupling to circular states [36], and optimal control techniques for improving state preparation efficiencies [37]. The long coherence times may also be used to extend simulations of quantum magnets to situations with strong spin-phonon coupling [38, 39, 40], and additional control of radio-frequency coupling to other high-$l$ states in the giant Rydberg manifold provides access to large-spin Heisenberg models [41].
@@ Appendix A Circularization
In this section we provide additional detail for the adiabatic circularization method described in the main text. By coupling the lowest lying states of each $m$, the initial optically accessible F-state can be coupled to the circular Rydberg state. To optimize the adiabatic multi-crossing efficiency when sweeping the electric field, the frequency of all involved transitions must be approximately equal $f_{m,m+1}\approx f_{\mathrm{RF}}$. Fig. 6(a) shows those frequencies, calculated (using [42]) with the quantum defects reported in [43] as a function of the electric field and magnetic field of $B=0.4\,\mathrm{G}$. The $m\leq 3$ states are split-off from the equally spaced hydrogenic manifold with $m\geq 4$ by the quantum defect. While the transition frequencies $f_{2,3}$ and $f_{3,4}$ attach smoothly to the transition frequency of the hydrogenic manifold $f_{\mathrm{HM}}$ with increasing electric field, the lower-$m$ states get further shifted by states of the neighboring $n=78$ manifold, decreasing the transition frequency below $f_{\mathrm{HM}}$. The electric field during the ARP must be chosen close to the point where the transition frequency from the initial $m=2$ state $f_{2,3}$ is close to $f_{\mathrm{HM}}$, while keeping $f_{1,2}$ out of resonance to prevent malicious population transfer to states with $m<2$. Note that the calculated crossing point is at $E=0.43\,\mathrm{Vcm^{-1}}$ whereas we find it to be between $E=0.46\,\mathrm{Vcm^{-1}}$ and $E=0.5\,\mathrm{Vcm^{-1}}$ in the experiment. We attribute this to uncertainties in the quantum defects. In Fig. 6(b), SSFI-traces after the $t_{\mathrm{ARP}}=20\,\mathrm{\upmu s}$ long ARP and a subsequent microwave transfer from |$\ket{79\mathrm{C}}$ to |$\ket{77\mathrm{C}}$ for varying RF voltage amplitudes applied to the circular electric field electrodes is shown. For low amplitudes $<6\,\mathrm{mV}$ the emergent avoided crossing is too weak and (partial) diabatic crossing to states with $m<n-1$ occurs. For amplitudes between $6\,\mathrm{mV}$ and $12\,\mathrm{mV}$ a clear, robust transfer to the |$\ket{77\mathrm{C}}$ is visible, indicating a high success rate of the ARP. If the amplitude is chosen even higher, the transfer efficiency suddenly drops drastically. This can be explained by a coupling to the $m=0,1$ states, if the split-off frequency $|f_{1,2}-f_{\mathrm{RF}}|$ from the $m\geq 4$ states gets comparable to the Rabi frequency of the RF-drive, leading to population transfer to those states.
@@ Appendix B Modeling Qubit Decoherence
Under the assumption that the noise is dominated by magnetic field noise we can also give an estimate of the magnetic field fluctuation magnitude
$\sigma_{B}=\frac{1}{T_{2}^{*}}\frac{1}{\delta_{m}\mu_{\mathrm{B}}},$ (9)
where $\mu_{\mathrm{B}}$ is the Bohr magneton and $\delta_{m}=2$ the difference in magnetic quantum number $m$ between the qubit states |$\ket{79\mathrm{C}}$ and |$\ket{77\mathrm{C}}$. With the data from Fig. 2 we find $\sigma_{B}\approx 1.33(7)\,\mathrm{mG}$, which is comparable with values we find from independent measurements of the microwave transition resonance.

### 04 Mar 2024 v2

@@ V Lifetime
Finally, we investigate the lifetime of our circular Rydberg qubit in more detail. To this end, we initialize the qubit either in |$\ket{79\mathrm{C}}$ or |$\ket{77\mathrm{C}}$, the latter via a microwave $\pi$-pulse applied after preparing |$\ket{79\mathrm{C}}$. For both scenarios, we hold the atoms in shallow tweezers ($\approx 400\,\mathrm{\upmu W}$) for a variable time $t$ and subsequently perform state-selective field ionization. This allows us to identify black-body induced population transfer into neighboring $n$-manifolds. Representative histograms of the field-ionization signal are depicted in Fig. 5(c) and (d) at $t\approx 630\,\mathrm{\upmu s}$ (indicated by dashed vertical lines) when starting in either of the qubit states. Evidently, decay from the circular state appears to be significantly slower for |$\ket{79\mathrm{C}}$. For a quantitative analysis, we fit the histograms for each hold time with a sum over multiple skewed Gaussians with the individual amplitudes as only free parameters and extract the time-dependent state population $p_{n}$ from the areas under the individual curves. Results are depicted in Fig. 5(a) and (b), and allow for extracting lifetimes $\tau_{\ket{n\mathrm{C}}}$ by fitting the data with a rate model (see Appendix E). Specifically, we obtain $\tau_{\ket{79\mathrm{C}}}=2.55(10)\,\mathrm{ms}$ and $\tau_{\ket{77\mathrm{C}}}=0.53(8)\,\mathrm{ms}$. Comparing these observations to predictions in free space at room temperature ($\tau_{\ket{79\mathrm{C}},\rm{fs}}=303\,\mathrm{\upmu s}$ and $\tau_{\ket{77\mathrm{C}},\rm{fs}}=297\,\mathrm{\upmu s}$), we find an enhancement in lifetime by factors of $8.4$ and $1.8$, respectively. The long circular state lifetimes are attributed to the presence of the ITO electrodes forming a plate capacitor in the $x$-$y$ plane, which coincides with the orbital plane of the Rydberg electron. The capacitor plates are spaced by $d=10.0(2)\,\mathrm{mm}$, which is slightly smaller than the half-wavelength of black-body transitions into the neighboring $n$-manifolds $\lambda_{\rm{BB}}^{n\rightarrow n-1}/2=10.6~{}(10.2)\,\mathrm{mm}$ for |$\ket{79\mathrm{C}}$ (|$\ket{77\mathrm{C}}$). This leads to suppression of the most detrimental circularly polarized black-body modes inside the electrode structure and allows us to create and control long-lived circular states without cryogenic cooling. The fact that the capacitor spacing is only slightly larger than the relevant black-body wavelength leads to the large difference in suppression factors we find for the two circular qubit states. We find good agreement with calculations for a infinite plate capacitor along the lines of [17], yielding $\tau_{\ket{77\mathrm{C}}}=0.56\,\mathrm{ms}$ and $\tau_{\ket{79\mathrm{C}}}=2.37\,\mathrm{ms}$ when assuming a capacitor spacing of $d=10.15\,\mathrm{mm}$ and plate reflectivity of $R=0.96$. As the lifetime is extremely sensitive to the capacitor spacing in this regime, measurements at higher $n$ and a simulation of our finite electrode geometry is necessary to fully characterize the suppression effect. Nevertheless, our calculations suggest that lifetimes $>$10 ms should be reached when increasing $n$ to $\geq 96$.
@@ VI Conclusion and Outlook
We have demonstrated creation and trapping of alkaline-earth atoms in very high-$n$ circular Rydberg states by applying radio-frequency driven adiabatic state preparation to a record-breaking principal quantum number of $n=79$. Trapping in a standard Gaussian tweezer beam is possible due to the core polarization of the divalent Rydberg atom. We implemented microwave control of a circular Rydberg atom qubit and characterized qubit coherence via Ramsey and spin-echo spectroscopy, including analysis of trap-induced dephasing. Our results show coherence times up to 278(10) $\upmu$s, limited by residual magnetic field noise and electric field gradients in our current setup. Assisted by an in-vacuum suppression capacitor for black-body modes, we find that the produced circular states live for up to 2.5 ms, which to the best of our knowledge is the longest-lived Rydberg atom ever observed in a room temperature environment.
Our work opens the door to quantum simulations with alkaline-earth circular Rydberg states, and provides prospects for novel qubit concepts in gate-based Rydberg quantum computers [13, 14, 17]. Higher preparation fidelities or partial-ionization methods to purify the created circular Rydberg states [36] will allow us to decrease the electric field we applied here to separate transitions of nearby elliptical states, which drastically decreases electric field sensitivity. Together with improved magnetic field noise cancellation, this should allow for reaching qubit coherence in the millisecond range and beyond [17], which provides exciting opportunities to overcome fundamental lifetime limitations in state-of-the-art quantum simulators using low-$l$ Rydberg states by orders of magnitude. Specifically, the high-$n$ circular states we demonstrate here give $>10\,\mathrm{MHz}$ exchange coupling between neighboring $n$, and thus on the order of $10^{4}-10^{5}$ coherent flip-flops within their lifetime even at room temperature. Perspectively, reaching two-body gate-fidelities beyond 99.9 % would also profit from increased Rydberg state lifetime, at the expense of additional challenges associated with using circular states for digital mode operation [14]. The available ionic-core transitions can be exploited in future experiments for optical read-out and laser cooling of circular Rydberg atoms [37, 38, 39]. Narrow-line spectroscopy on the ${}^{+}$ core also enables coherent and local optical manipulation of the circular state qubit mediated by the electrostatic coupling between the two electrons [27].
Next obvious steps include the creation and control of interacting circular Rydberg atoms, scaling the system to large qubit arrays, possibly assisted by rapid Rabi coupling to circular states [40], and optimal control techniques for improving state preparation efficiencies [41, 42]. Moreover, dynamical rearrangement of the Rydberg atoms within their long lifetime is in reach, and together with non-destructive state detection [27], e.g. via low-$l$ ancilla atoms [14], could be used for filling defects in the array. The long coherence times may then be used to extend simulations of quantum magnets to situations with strong spin-phonon coupling [43, 44, 45], and additional control of radio-frequency coupling to other high-$l$ states in the giant Rydberg manifold provides access to large-spin Heisenberg models [46].
@@ Appendix A Circularization
In this section we provide additional detail for the adiabatic circularization method described in the main text. By coupling the lowest lying states of each $m$, the initial optically accessible F-state can be coupled to the circular Rydberg state. To optimize the adiabatic multi-crossing efficiency when sweeping the electric field, the frequency of all involved transitions must be approximately equal $f_{m,m+1}\approx f_{\mathrm{RF}}$. Fig. 6(a) shows those frequencies, calculated (using [47]) with the quantum defects reported in [48] as a function of the electric field and magnetic field of $B=0.4\,\mathrm{G}$. The $m\leq 3$ states are split-off from the equally spaced hydrogenic manifold with $m\geq 4$ by the quantum defect. While the transition frequencies $f_{2,3}$ and $f_{3,4}$ attach smoothly to the transition frequency of the hydrogenic manifold $f_{\mathrm{HM}}$ with increasing electric field, the lower-$m$ states get further shifted by states of the neighboring $n=78$ manifold, decreasing the transition frequency below $f_{\mathrm{HM}}$. The electric field during the ARP must be chosen close to the point where the transition frequency from the initial $m=2$ state $f_{2,3}$ is close to $f_{\mathrm{HM}}$, while keeping $f_{1,2}$ out of resonance to prevent malicious population transfer to states with $m<2$. Note that the calculated crossing point is at $E=0.43\,\mathrm{Vcm^{-1}}$ whereas we find it to be between $E=0.46\,\mathrm{Vcm^{-1}}$ and $E=0.5\,\mathrm{Vcm^{-1}}$ in the experiment. We attribute this to uncertainties in the quantum defects which are measured at $n=50$. In Fig. 6(b), SSFI-traces after the $t_{\mathrm{ARP}}=20\,\mathrm{\upmu s}$ long ARP and a subsequent microwave transfer from |$\ket{79\mathrm{C}}$ to |$\ket{77\mathrm{C}}$ for varying RF voltage amplitudes applied to the circular electric field electrodes is shown. For low amplitudes $<6\,\mathrm{mV}$ the emergent avoided crossing is too weak and (partial) diabatic crossing to states with $m<n-1$ occurs. For amplitudes between $6\,\mathrm{mV}$ and $12\,\mathrm{mV}$ a clear, robust transfer to the |$\ket{77\mathrm{C}}$ is visible, indicating a high success rate of the ARP. If the amplitude is chosen even higher, the transfer efficiency suddenly drops drastically. This can be explained by a coupling to the $m=0,1$ states, if the split-off frequency $|f_{1,2}-f_{\mathrm{RF}}|$ from the $m\geq 4$ states gets comparable to the Rabi frequency of the RF-drive, leading to population transfer to those states.
@@
Since the circularization parameters are very sensitive to the F-state attaching to the manifold, the quantum defects are likely not accurate enough to find the optimal parameters. We therefore propose that the preparation fidelity can be strongly increased from the currently 70 % by an in-depth optimization, preferably guided by optimal control methods. Since the parameter space including all electric, magnetic and radio frequency fields is very high-dimensional we leave this for future work. Perspectively, we expect to reach also preparation times on the order of 100 ns and fidelities of 99 % by optimal control assisted coherent excitation as reported in [42] for lower $n$.
@@ Appendix B Modeling Qubit Decoherence
With the qubit transition sensitive to magnetic and electric fields, we identify four possible decoherence sources to the $T_{2}^{*}$ time: temporal fluctuations and spatial gradients, both in the electric and magnetic field. From independent Rydberg F-state spectroscopy at varying electric fields in single tweezers, we find an upper bound for electric field fluctuations of $\delta_{E}=50\,\mathrm{\upmu V/cm}$. With a sensitivity of the |$\ket{79\mathrm{C}}\rightarrow\ket{77\mathrm{C}}$ transition of $\xi_{E}=2\pi\times 8.8\,\mathrm{kHz(mV/cm)^{-1}}$ at the electric control field of $E=2\,\mathrm{V/cm}$ used throughout this work, we find a variance of the effective noise-induced detuning of
$\sigma_{E}=\xi_{E}\delta_{E}\approx 2\pi\times 0.44\,\mathrm{kHz}.$ (9)
From similar spectroscopy measurements, we find that the electric field gradient across the tweezer array is $\Delta_{E}=1.1\,\mathrm{mV/mm^{2}}$. The gradient causes a discrete, constant detuning on each tweezer, for which the variance can be obtained from the sum over all tweezer sites $k$
$\sigma_{\Delta_{E}}^{2}=\frac{1}{10}\sum_{k=0}^{9}\left(kd\Delta_{E}\xi_{E}-\mu_{E}\right)^{2}=\left(2\pi\times 2.8\,\mathrm{kHz}\right)^{2},$ (10)
with the mean $\mu_{E}=\frac{1}{10}\sum_{k=0}^{9}\left(kd\Delta_{E}\xi_{E}\right)$ and the distance between tweezers $d=10\,\mathrm{\upmu m}$. Analogously, we obtain $\sigma_{\Delta_{B}}=2\pi\times 20\,\mathrm{mHz}$ for a magnetic field gradient of $\Delta_{B}=26\,\mathrm{mG/cm}$ we find in our experiment and a magnetic sensitivity of the transition of $\xi_{B}=2\pi\times\delta_{m}\mu_{B}$ with $\delta_{m}=2$ the difference in magnetic quantum number $m$, and $\mu_{B}=1.4\,\mathrm{MHz/G}$. Evidently, decoherence by magnetic field gradients and electric field fluctuations are orders of magnitude smaller than the influence of the electric field gradient and can be neglected. Assuming uncorrelated noise sources, the $T_{2}^{*}$ time is directly connected to the summed variances via
$\frac{1}{{T_{2}^{*}}^{2}}=\sigma_{\Delta E}^{2}+\sigma_{B}^{2}.$ (11)
With the measured $T_{2}^{*}$ time we can give an approximate value for the magnetic field noise by solving Eq. 11 for $\sigma_{B}$. We find $\sigma_{B}=2\pi\times 2.4\,\mathrm{kHz}$ on the same order as the electric field gradient induced noise. This corresponds to magnetic field fluctuations of $\delta_{B}\approx 0.9\,\mathrm{mG}$, comparable with values we find from independent measurements of the microwave transition resonance.